Della takes 12 hours to dust the house. It takes Bebe two hours to finish this job. How long would it take if they worked together? <-- *can you make the representation using x thankyou*
Answer by CubeyThePenguin(3113)

Della: 1/12 of the house dusted in 1 hr
Bebe: 1/2 of the house dusted in 1 hr

1/12 + 1/2 = 14/24 = 7/12

It takes 12/7 hours for both of them to finish.